# Large Language Model-Brained GUI Agents: A Survey

### v9

@@ 2.2 Surveys on LLM Agents
Our survey distinguishes itself from existing work by providing a pioneering and comprehensive analysis of the intersections and integrations between LLMs and GUI agents. It thoroughly examines foundational components, advanced technologies, framework architectures, data and model optimization strategies, applications, key challenges, and a forward-looking roadmap—areas that have not been fully explored in previous studies.
@@ 4.3.4 Industry Models
Each type has distinct advantages, and a deeper understanding of these approaches is critical for designing effective agents. Specifics of these action spaces are discussed in Section 5.5.
@@
The most effective digital agents are likely to operate in a hybrid manner, combining the strengths of both approaches. Such agents can utilize GUI operations to achieve broad compatibility across software while exploiting native APIs where available to maximize efficiency and effectiveness. These hybrid agents strike a balance between generalization and task optimization, making them a critical focus area in this survey.
@@ 5.2.1 Platform
GUI agents can interact with a wide range of platforms, including mobile devices, web applications, and computer operating systems like Windows. Each platform offers unique capabilities and constraints for GUI automation, requiring agents to adapt their perception and interaction strategies accordingly.
@@ 5.2.2 Environment State Perception
TABLE III: Key toolkits for collecting GUI environment data.
@@
Accurately perceiving the current state of the environment is essential for LLM-powered GUI agents, as it directly informs their decision-making and action-planning processes. This perception is enabled by gathering a combination of structured data, such as widget trees, and unstructured data, like screenshots, to capture a complete representation of the interface and its components. In Table III, we outline key toolkits available for collecting GUI environment data across various platforms, and below we discuss their roles in detail:
@@ 5.3 Prompt Engineering
    Agent Instruction: This section provides guidance for the agent’s operation, detailing its role, rules to follow, and specific objectives. Instructions clarify what inputs the agent will receive and outline the expected outputs from the LLM, establishing a framework for the inference process.
@@ 5.4.1 Planning
Successful execution of GUI tasks often necessitates a series of sequential actions, requiring the agent to engage in effective planning [183]. Analogous to human cognitive processes, thoughtful planning is essential to organize tasks, schedule actions, and ensure successful completion [51, 184]. The LLM must initially conceptualize a long-term goal while simultaneously focusing on short-term actions to initiate progress toward that goal [185].
@@
TABLE IV: Overview of actions for GUI agents.
@@ 5.4.3 Complementary Outputs
In addition to planning and action inference, the LLM can also generate complementary outputs that enhance the agent’s capabilities. These outputs may include reasoning processes that clarify the agent’s decision-making (e.g., CoT reasoning), messages for user interaction, or communication with other agents or systems, or the status of the task (e.g., continue or finished). The design of these functionalities can be tailored to meet specific needs, thereby enriching the overall performance of the GUI agent.
@@ 5.5 Actions Execution
Following the inference process, a crucial next step is for the GUI agent to execute the actions derived from the inferred commands within the GUI environment and subsequently gather feedback. Although the term “GUI agent” might suggest a focus solely on user interface actions, the action space can be greatly expanded by incorporating various toolboxes that enhance the agent’s versatility. Broadly, the actions available to GUI agents fall into three main categories: (i) UI operations [138], (ii) native API calls [190], and (iii) AI tools [191]. Each category offers unique advantages and challenges, enabling the agent to tackle a diverse range of tasks more effectively. We summarize the various actions commonly used in GUI agents, categorized into distinct types, in Table IV, and provide detailed explanations of each category below.
@@ 5.6 Memory
TABLE V: Summary of memory in GUI agents.
@@
For a GUI agent to achieve robust performance in complex, multi-step tasks, it must retain memory, enabling it to manage states in otherwise stateless environments. Memory allows the agent to track its prior actions, their outcomes, and the task’s overall status, all of which are crucial for informed decision-making in subsequent steps [194]. By establishing continuity, memory transforms the agent from a reactive system into a proactive, stateful one, capable of self-adjustment based on accumulated knowledge. The agent’s memory is generally divided into two main types: Short-Term Memory [195] and Long-Term Memory [196]. We show an overview of different types of memory in GUI agents in Table V.
@@ 5.6.2 Long-Term Memory
LTM contributes to the agent’s self-improvement over time by preserving examples of successful task trajectories, operational guidelines, and common interaction patterns. When approaching a new task, the agent can leverage RAG techniques to retrieve relevant historical data, which enhances its ability to adapt strategies based on prior success. This makes LTM instrumental in fostering an agent’s capacity to “learn” from experience, enabling it to perform tasks with greater accuracy and efficiency as it accumulates insights across sessions. For instance, [199] provides an illustrative example of using past task trajectories stored in memory to guide and enhance future decision-making, a technique that is highly adaptable for GUI agents. It also enables better personalization by retaining information about previous tasks.
@@ 5.7 Advanced Enhancements
TABLE VI: A summary of of GUI grounding models and benchmarks.
@@
ScreenSpot [25] iOS, Android, macOS, and Windows (benchmark) (benchmark) Over 600 screenshots and 1,200 instructions GUI screenshots accompanied by user instructions Bounding boxes or coordinates of actionable GUI elements A realistic and diverse GUI grounding benchmark covering multiple platforms and a variety of elements https://github.com/njucckevin/SeeClick
ScreenSpot-Pro [220] Windows, macOS, and Linux (benchmark) (benchmark) 1,581 instruction–screenshot pairs covering 23 applications across 5 industries and 3 operating systems High-resolution GUI screenshots paired with natural language instructions Bounding boxes for locating target UI elements Introduces a high-resolution benchmark for professional environments https://github.com/likaixin2000/ScreenSpot-Pro-GUI-Grounding
@@ 5.7.1 Computer Vision-Based GUI Grounding
Although various tools (Section III) enable GUI agents to access information like widget location, captions, and properties, certain non-standard GUIs or widgets may not adhere to these tools’ protocols [221], rendering their information inaccessible. Additionally, due to permission management, these tools are not always usable. Such incomplete information can present significant challenges for GUI agents, as the LLM may need to independently locate and interact with required widgets by estimating their coordinates to perform actions like clicking—a task that is inherently difficult without precise GUI data.
CV models offer a non-intrusive solution for GUI grounding directly from screenshots, enabling the detection, localization, segmentation, and even functional estimation of widgets [222, 97, 223, 224]. This approach allows agents to interpret the visual structure and elements of the GUI without relying on system-level tools or internal metadata, which may be unavailable or incomplete. CV-based GUI parsing provides agents with valuable insights into interactive components, screen layout, and widget functionalities based solely on visual cues, enhancing their ability to recognize and act upon elements on the screen. Figure 9 provides an illustrative example of how a CV-based GUI parser works. While standard API-based detection captures predefined widgets, the CV model can identify additional elements, such as thumbnails and canvases, which may not have explicit API representations in the PowerPoint interface. This enhances widget recognition, allowing the agent to detect components beyond the scope of API detection. We show an overview of related GUI grounding models and benchmarks in Table VI.
@@ 5.7.4 Self-Evolution
    Guidance and Rules: From the accumulated experiences, agents can extract high-level rules or guidelines that encapsulate best practices, successful strategies, and lessons learned from past mistakes [234, 235]. This knowledge can be formalized into policies or heuristics that agents consult during decision-making processes, thereby enhancing their reasoning capabilities.
@@
    New Toolkits: Throughout their interactions, GUI agents may discover or develop more efficient methods, tools, or sequences of actions that streamline task execution [155]. These may include optimized API calls, macros, or combinations of UI operations that accomplish tasks more effectively than previous approaches. By incorporating these new tools into their repertoire, agents expand their capabilities and enhance overall efficiency.
@@ 5.7.5 Reinforcement Learning
Reinforcement Learning (RL) [236] has witnessed significant advancements in aligning LLMs with desired behaviors [237], and has recently been employed in the development of LLM agents [50, 238, 238]. In the context of GUI multi-agent systems, RL offers substantial potential to enhance the performance, adaptability, and collaboration of GUI agents. GUI automation tasks naturally align with the structure of a Markov Decision Process (MDP) [239], making them particularly well-suited for solutions based on RL. In this context, the state corresponds to the environment perception (such as GUI screenshots, UI element properties, and layout configurations), while actions map directly to UI operations, including mouse clicks, keyboard inputs, and API calls. Rewards can be explicitly defined based on various performance metrics, such as task completion, efficiency, and accuracy, allowing the agent to optimize its actions for maximal effectiveness. Figure 17 illustrates an example of MDP modeling for task completion in a GUI agent, where state, action and reward are clearly defined.
@@ 5.8 From Foundations to Innovations: A Roadmap
FoundationstoInnovations Frameworks Web [243] [17, 193] [244, 245, 246, 247, 248, 249, 250, 251, 252, 253, 254, 255, 256, 257, 258, 259, 260, 261, 262, 263, 264, 265, 266, 267, 268, 269, 270, 271] Mobile [18, 227, 152] [150, 272, 266] [260, 261, 267] [270, 273, 274, 275, 276, 277, 278, 279, 280, 281, 282, 283, 284, 285, 286, 287, 288, 289, 290] Computer [270, 243] [19, 155, 156] [227, 278, 261] [291, 292, 293, 284] [266, 267, 294] Cross- Platform [243] [260, 266] [261, 227] [284, 295] [267, 270] Data Web [206, 211] [296, 297, 298, 299, 300, 301, 302, 303, 304, 305] Mobile [177, 138] [272, 298] [211, 299] [303, 304] [306, 307, 308, 309, 310, 311, 312, 313, 314, 315, 316, 317, 318, 319, 320] Computer [177, 303] [321, 319, 320, 322] [305] Cross- Platform [304, 305] [319, 303] [177, 211] [299, 301] [298, 320] Models Foundation Models [159] [200, 204] [86, 87, 157] [323, 324, 325, 326, 327, 328, 329, 330, 331] Web [25, 177, 332] [147, 259, 301] [303, 333, 334, 335, 336, 337, 338, 339, 340] Mobile [25, 15, 177] [241, 306, 316] [314, 341, 313, 342, 343, 344, 345, 153, 346, 347, 320, 338, 339, 340] [303, 333, 348] Computer [25, 15, 321] [349, 347, 320] [177, 339, 303] [340, 301, 322] [333, 332] Cross- Platform [15, 25, 303] [177, 320, 301] [347, 338, 339, 340] [333, 332] Evaluation Web [207] [256, 248, 25] [136, 139, 140] [206, 296, 297, 298] [350, 351, 352, 353, 300, 354, 355, 356, 357, 358, 359, 360, 361, 362, 363, 364, 365, 366, 367, 368] Mobile [317, 365] [369, 150, 138] [240, 152, 370, 371, 372, 310, 373, 374, 375, 376, 298, 25, 377, 378, 379, 380, 381] Computer [25, 369] [365, 103, 382] [357, 358, 359] [383, 384, 385, 386] Cross- Platform [292] [298, 358, 359] [369, 25, 365] Applications GUI Testing [387, 119] [388, 389, 390, 391, 392, 393, 394, 395, 396, 397] Virtual Assistants [126] [398, 399, 400, 401, 402, 403, 404, 405, 406, 407, 408, 409, 410]

### v11

@@ 2 Related Work
Aghzal et al., [52] A survey on automated planning in LLMs.
Zheng et al., [53] Discuss the roadmap of lifelong learning in LLM agents.
@@
Shi et al., [67] A survey of building trustworthy GUI agents.
Ning et al., [68] A survey of agents for Web automation.
Tang et al., [69] A survey of GUI agents powered by (multimodal) LLMs.
@@ 2.2 Surveys on LLM Agents
Compared to existing surveys, our work offers a significantly more comprehensive and up-to-date overview of the LLM-powered GUI agent landscape. We curate and synthesize over 500 references, covering a wide range of topics including foundation models, data sources, system frameworks, benchmarks, evaluation methodologies, and practical deployments. While prior surveys often concentrate on narrower aspects on selected platform (e.g., web, mobile), our survey takes a holistic perspective that spans the full development and deployment lifecycle. Beyond narrative summaries, we also provide consolidated reference tables for each subdomain, enabling readers to quickly categorize and locate relevant works across platforms and research themes—serving as a practical handbook for both researchers and practitioners. Furthermore, we incorporate foundational background material and propose evaluation taxonomies that make the survey accessible to newcomers, addressing gaps in prior work that often assume a high degree of prior familiarity.
@@ 4.3.4 Industry Models
We show the principle of both agent types in Figure 4. Each type has distinct advantages, and a deeper understanding of these approaches is critical for designing effective agents.
@@
The most effective digital agents are likely to operate in a hybrid manner, combining the strengths of both approaches. Such agents can utilize GUI operations to achieve broad compatibility across software while exploiting native APIs where available to maximize efficiency and effectiveness. These hybrid agents strike a balance between generalization and task optimization, making them a critical focus area in this survey. For a more comprehensive comparison between GUI agents and API agents, please refer to [165].
@@ 5.1 Architecture and Workflow In a Nutshell
TABLE III: Summary of platform-specific challenges, action spaces, and typical tasks for Web, Mobile, and Computer GUI environments.
Platform Typical GUI Challenges Action Space Representative Tasks
Mobile Constrained screen real estate Heavy reliance on touch and gesture recognition [168] App architectures (native vs. hybrid) Accessibility frameworks (e.g., Android’s Accessibility API, iOS VoiceOver) Platform-specific constraints (permissions, security, privacy) Tap, swipe, pinch, and other touch gestures Virtual keyboard input In-app navigation (menus, tabs) Accessing hardware features (camera, GPS) App-based login and form filling Messaging, social media posting Location-based services and map interactions Handling push notifications and permission dialogs
Web Dynamic and responsive layouts Asynchronous updates (AJAX, fetch APIs) HTML/DOM-based structures Cross-browser inconsistencies Click, hover, scroll DOM-based form filling Link navigation and element inspection JavaScript event triggering Form completion (registrations, checkouts) Data extraction/web scraping Searching and filtering (e.g., e-commerce) Multi-step web navigation (redirects, pop-ups)
Computer Full-fledged OS-level interfaces Multi-window operations and system-level shortcuts Automation APIs (e.g., Windows UI Automation [32]) Frequent software updates requiring adaptation Complex, multi-layered software suites Mouse click, drag-and-drop Keyboard shortcuts and text input Menu navigation, toolbars Access to multiple application windows File management and system settings Productivity software usage (office suites, IDEs) Installing/uninstalling applications Coordinating multi-application workflows
@@ 5.2.2 Environment State Perception
TABLE IV: Key toolkits for collecting GUI environment data.
@@
Accurately perceiving the current state of the environment is essential for LLM-powered GUI agents, as it directly informs their decision-making and action-planning processes. This perception is enabled by gathering a combination of structured data, such as widget trees, and unstructured data, like screenshots, to capture a complete representation of the interface and its components. In Table IV, we outline key toolkits available for collecting GUI environment data across various platforms, and below we discuss their roles in detail:
@@ 5.3 Prompt Engineering
    Agent Instruction: This section provides guidance for the agent’s operation, detailing its role, rules to follow, and specific objectives. Instructions clarify what inputs the agent will receive and outline the expected outputs from the LLM, establishing a framework for the inference process. The core agent instructions are usually embedded within the base system prompt of the LLM, with supplementary instructions dynamically injected or updated based on environmental feedback and contextual adaptation.
@@ 5.4.1 Planning
Successful execution of GUI tasks often necessitates a series of sequential actions, requiring the agent to engage in effective planning [189, 52]. Analogous to human cognitive processes, thoughtful planning is essential to organize tasks, schedule actions, and ensure successful completion [51, 190]. The LLM must initially conceptualize a long-term goal while simultaneously focusing on short-term actions to initiate progress toward that goal [191].
@@
TABLE V: Overview of actions for GUI agents.
@@ 5.4.3 Complementary Outputs
In addition to planning and action inference, the LLM can also generate complementary outputs that enhance the agent’s capabilities. These outputs may include reasoning processes that clarify the agent’s decision-making (e.g., CoT reasoning [196]), messages for user interaction, or communication with other agents or systems, or the status of the task (e.g., continue or finished). The design of these functionalities can be tailored to meet specific needs, thereby enriching the overall performance of the GUI agent.
@@ 5.5 Actions Execution
Following the inference process, a crucial next step is for the GUI agent to execute the actions derived from the inferred commands within the GUI environment and subsequently gather feedback. Although the term “GUI agent” might suggest a focus solely on user interface actions, the action space can be greatly expanded by incorporating various toolboxes that enhance the agent’s versatility. Broadly, the actions available to GUI agents fall into three main categories: (i) UI operations [143], (ii) native API calls [197], and (iii) AI tools [198]. Each category offers unique advantages and challenges, enabling the agent to tackle a diverse range of tasks more effectively. We summarize the various actions commonly used in GUI agents, categorized into distinct types, in Table V, and provide detailed explanations of each category below.
@@ 5.6 Memory
TABLE VI: Summary of memory in GUI agents.
@@
For a GUI agent to achieve robust performance in complex, multi-step tasks, it must retain memory, enabling it to manage states in otherwise stateless environments. Memory allows the agent to track its prior actions, their outcomes, and the task’s overall status, all of which are crucial for informed decision-making in subsequent steps [201]. By establishing continuity, memory transforms the agent from a reactive system into a proactive, stateful one, capable of self-adjustment based on accumulated knowledge. The agent’s memory is generally divided into two main types: Short-Term Memory [202] and Long-Term Memory [203]. We show an overview of different types of memory in GUI agents in Table VI.
@@ 5.6.2 Long-Term Memory
LTM contributes to the agent’s self-improvement over time by preserving examples of successful task trajectories, operational guidelines, and common interaction patterns. When approaching a new task, the agent can leverage RAG techniques to retrieve relevant historical data, which enhances its ability to adapt strategies based on prior success. This is similar to the lifelong learning [53], which makes LTM instrumental in fostering an agent’s capacity to “learn” from experience, enabling it to perform tasks with greater accuracy and efficiency as it accumulates insights across sessions. For instance, [206] provides an illustrative example of using past task trajectories stored in memory to guide and enhance future decision-making, a technique that is highly adaptable for GUI agents. It also enables better personalization by retaining information about previous tasks.
@@ 5.7 Advanced Enhancements
TABLE VII: A summary of of GUI grounding models (Part I).
@@
UGround [217] Web, Desktop (Windows, MacOS, Linux), Mobile (Android, iOS) LLaVA-NeXT-7B [218] 7B Web-Hybrid and other existing datasets GUI screenshots, user queries Pixel coordinates of GUI elements A universal GUI grounding model that relies solely on vision, eliminating the need for text-based representations https://osu-nlp-group.github.io/UGround/
@@
TABLE VIII: A summary of of GUI grounding models (Part II).
@@
AutoGUI [229] Web, Mobile Qwen-VL-10B [211], SliME-8B [230] 10B / 8B AutoGUI-704k GUI screenshots, User queries Element functionalities, Element locations Automatically labels UI elements based on interaction-induced changes, making it scalable and high-quality. https://autogui-project.github.io/
Query Inference [231] Mobile Android Qwen2-VL-7B-Instruct [232] 7B UIBERT [233] GUI screenshots Action-oriented queries, Coordinates Improves reasoning without requiring large-scale training data. https://github.com/ZrW00/GUIPivot
WinClick [234] Windows OS Phi3-Vision [235] 4.2B WinSpot Benchmark GUI screenshots, Natural language instructions Element locations The first GUI grounding model specifically tailored for Windows. https://github.com/zackhuiiiii/WinSpot
FOCUS [236] Web, mobile applications, and desktop Qwen2-VL-2B-Instruct [232] 2B GUICourse [220], Aguvis-stage1 [237], Wave-UI [238], Desktop-UI [239] GUI screenshot + task instruction Normalized coordinates (x, y) A dual-system GUI grounding architecture inspired by human cognition, which dynamically switches between fast (intuitive) and slow (analytical) grounding modes based on task complexity https://github.com/sugarandgugu/Focus
UI-E2I-Synth [240] Web, Windows, and Android InternVL2-4B and Qwen2-VL-7B 4B and 7B 1.6M screenshots, 9.9M instructions GUI screenshot Element coordinates Introduces a three-stage synthetic data pipeline for GUI grounding with both explicit and implicit instruction synthesis https://colmon46.github.io/i2e-bench-leaderboard/
TABLE IX: A summary of of GUI grounding benchmarks.
Benchmark Platform Dataset Input Output Highlight Link
ScreenSpot [25] iOS, Android, macOS, and Windows Over 600 screenshots and 1,200 instructions GUI screenshots accompanied by user instructions Bounding boxes or coordinates of actionable GUI elements A realistic and diverse GUI grounding benchmark covering multiple platforms and a variety of elements https://github.com/njucckevin/SeeClick
ScreenSpot-Pro [241] Windows, macOS, and Linux 1,581 instruction–screenshot pairs covering 23 applications across 5 industries and 3 operating systems High-resolution GUI screenshots paired with natural language instructions Bounding boxes for locating target UI elements Introduces a high-resolution benchmark for professional environments https://github.com/likaixin2000/ScreenSpot-Pro-GUI-Grounding
PixelWeb [242] Web 100,000 webpages Rendered webpage screenshots and DOM information BBox, mask, contour The first GUI dataset to provide pixel-level annotations—including mask and contour—for web UIs, enabling high-precision GUI grounding and detection tasks https://huggingface.co/datasets/cyberalchemist/PixelWeb
@@ 5.7.1 Computer Vision-Based GUI Grounding
Although various tools (Section IV) enable GUI agents to access information like widget location, captions, and properties, certain non-standard GUIs or widgets may not adhere to these tools’ protocols [243], rendering their information inaccessible. Additionally, due to permission management, these tools are not always usable. Such incomplete information can present significant challenges for GUI agents, as the LLM may need to independently locate and interact with required widgets by estimating their coordinates to perform actions like clicking—a task that is inherently difficult without precise GUI data.
CV models offer a non-intrusive solution for GUI grounding directly from screenshots, enabling the detection, localization, segmentation, and even functional estimation of widgets [244, 102, 245, 246]. This approach allows agents to interpret the visual structure and elements of the GUI without relying on system-level tools or internal metadata, which may be unavailable or incomplete. CV-based GUI parsing provides agents with valuable insights into interactive components, screen layout, and widget functionalities based solely on visual cues, enhancing their ability to recognize and act upon elements on the screen. Figure 10 provides an illustrative example of how a CV-based GUI parser works. While standard API-based detection captures predefined widgets, the CV model can identify additional elements, such as thumbnails and canvases, which may not have explicit API representations in the PowerPoint interface. This enhances widget recognition, allowing the agent to detect components beyond the scope of API detection. We show an overview of related GUI grounding models and benchmarks in Table VII, VIII and IX.
@@ 5.7.4 Self-Evolution
    Guidance and Rules: From the accumulated experiences, agents can extract high-level rules or guidelines that encapsulate best practices, successful strategies, and lessons learned from past mistakes [256, 257]. Such guidance can be acquired by the LLM itself through trajectory summarization [256], or even via search-based algorithms, such as Monte Carlo Tree Search (MCTS) [257]. This knowledge can be formalized into policies or heuristics that agents consult during decision-making processes, thereby enhancing their reasoning capabilities.
@@
    New Toolkits: Throughout their interactions, GUI agents may discover or develop more efficient methods, tools, or sequences of actions that streamline task execution [160]. These may include optimized API calls, macros, or combinations of UI operations that accomplish tasks more effectively than previous approaches. LLMs can be leveraged to automatically analyze execution trajectories in order to summarize, discover, and generate high-level shortcuts or frequently used fast APIs, which can then be reused for future executions [258]. By incorporating these new tools into their repertoire, agents expand their capabilities and enhance overall efficiency.
@@ 5.7.5 Reinforcement Learning
Reinforcement Learning (RL) [259] has witnessed significant advancements in aligning LLMs with desired behaviors [260], and has recently been employed in the development of LLM agents [50, 261]. In the context of GUI multi-agent systems, RL offers substantial potential to enhance the performance, adaptability, and collaboration of GUI agents. GUI automation tasks naturally align with the structure of a Markov Decision Process (MDP) [262], making them particularly well-suited for solutions based on RL. In this context, the state corresponds to the environment perception (such as GUI screenshots, UI element properties, and layout configurations), while actions map directly to UI operations, including mouse clicks, keyboard inputs, and API calls. Rewards can be explicitly defined based on various performance metrics, such as task completion, efficiency, and accuracy, allowing the agent to optimize its actions for maximal effectiveness. Figure 18 illustrates an example of MDP modeling for task completion in a GUI agent, where state, action and reward are clearly defined.
@@ 5.8 From Foundations to Innovations: A Roadmap
FoundationstoInnovations Frameworks Web [17, 200] [266, 267, 268, 269, 270, 271, 272, 273, 274, 275, 276, 277, 278, 279, 280, 281, 282, 283, 284, 285, 286, 287, 288, 289, 290, 291, 292, 293, 294, 295, 296, 297] Mobile [18, 157] [155, 298] [299, 300, 301, 302, 303, 304, 305, 306, 307, 308, 309, 310, 311, 312, 313, 314, 315, 316, 317, 258, 318, 319, 320, 321] Computer [19, 160] [161, 304] [322, 323, 324, 325, 326, 327, 328, 329, 330] Cross- Platform [331, 237] [332, 249] [333, 334, 335, 336, 337, 338, 339, 340] Data Web [213] [341, 342, 343, 344, 345] Mobile [143, 298] [346, 347, 348, 349, 350, 351, 352, 353, 354, 355, 356, 357, 358, 359] Computer [360, 361, 362] Cross- Platform [183] [220, 217] [363, 364, 365, 366, 367, 368, 369, 370, 371] Models Foundation Models [232, 235] [164, 207, 211] [91, 92, 162] [372, 373, 374, 375, 376, 218, 377, 378] Web [152, 281] [379, 380, 381, 382, 383] Mobile [355, 264] [346, 347, 348] [384, 385, 386, 387, 388, 158, 389, 390, 391, 392, 393, 394, 395, 383, 396] Computer [360, 397] [361, 398] Cross- Platform [15] [25, 367, 399] [183, 369, 366] [400, 239, 233] [401, 402, 403, 378, 404, 405] Evaluation Web [214] [278, 270] [141, 144, 145] [213, 341, 342] [406, 407, 408, 409, 343, 410, 411, 412, 413, 414, 415, 416, 417, 418, 419, 420, 421, 422, 423, 424, 425, 426, 427, 428, 429, 430] Mobile [352] [357, 155, 143] [263, 157, 431, 432, 433, 434, 435, 436, 437, 438, 439, 440, 441, 442, 443, 444, 445] Computer [108, 446, 413] [447, 448, 238, 449, 450, 451] Cross- Platform [323] [368, 452, 453] [454, 25, 455] Applications GUI Testing [456, 124] [457, 458, 459, 460, 461, 462, 463, 464, 465, 466, 467, 468, 469, 470, 471, 472, 473, 474] Virtual Assistants [131] [475, 476, 477, 478, 479, 480, 481, 482, 483, 484, 485, 486, 487, 488]